OpenStudy (anonymous):

If tan x=z/10 and cos x=10/y, what is the value of the sinx?

OpenStudy (anonymous):



OpenStudy (anonymous):

This triangle will be confusing :3

OpenStudy (anonymous):

sin(x) is just cos(x) times tan(x) ^^

OpenStudy (anonymous):

You know... coz \[\Large \tan(x) = \frac{\sin(x)}{\cos(x)}\]

OpenStudy (cj49):

sinθ=opp/hyp=z/y cosθ=adj/hyp=10/y tanθ=opp/adj=z/10

OpenStudy (kainui):

What does SOH CAH TOA mean? @Samidancer16

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

I think I already bluntly gave you all but the answer in my second comment... you should read it again ^^

OpenStudy (anonymous):

Can you say the answer again because i dont think they have that for an answer choice

OpenStudy (anonymous):

Of course it isn't... it's more of a strong hint :) Giving direct answers isn't allowed on here...

OpenStudy (cj49):

tanx = [sinx/cosx] z/10=sin÷10/y

OpenStudy (anonymous):

the answer choices are a) sinx=z/y b) sin x =y/z c)nsinx = 10z and d) sin x =10y

OpenStudy (anonymous):

Well, it's gotta be one of those. What happens when you multiply cos(x) with tan(x)?

OpenStudy (precal):